Find the dual program of the following linear programming problem.

π‘€π‘Žπ‘₯π‘–π‘šπ‘–π‘§π‘’ 𝑧 = 5π‘₯1 βˆ’ 2π‘₯2

𝑠𝑒𝑏𝑗𝑒𝑐𝑑 π‘‘π‘œ

3π‘₯1 + 2π‘₯2 β‰₯ 16

π‘₯1 βˆ’ π‘₯2 ≀ 4

π‘₯1 β‰₯ 5

π‘₯1 β‰₯0,π‘₯2 𝑖𝑠 π‘’π‘›π‘π‘œπ‘›π‘ π‘‘π‘Ÿπ‘Žπ‘–π‘›π‘’π‘‘Β 


1
Expert's answer
2021-09-27T03:06:57-0400

All constraints can be converted to "\leq" by multiplying by -1. So we have;

"\text{Max} ~~z=5x_1-2x_2\\
\text{subject to}\\
~~~~~-3x_1-2x_2\leq-16\\
~~~~~~~~~~~x_1-~~x_2\leq4\\
~~~~~~~-x_1~~~~~~~~~~~\leq-5\\
\text{and } x_1\leq 0; x_2 \text{is unrestricted}."

Since the primal has two variables and three constraints, then the dual will have three variables and two constraints. Also, the "x_2" variable is unrestricted in the primal, therefore the second constraint in the dual shall be equality.

Dual program is;

"\text{Min } ~~z=-16y_1+4y_2-5y_3\\
\text{subject to }\\
~~~~~-3y_1+y_2-y_3\geq5\\
~~~~~-2y_1-y_2~~~~~~~~=-2\\
\text{and } y_1,y_2\geq0."